OpenStudy (anonymous):

Solve for x: 2|2x - 2| + 4 = 20. x = 5, x = -3 x = 5, x = -5 x = 3, x = -3 x = -5, x = 3

OpenStudy (nurali):

2|2x - 2| + 4 = 20 2|2x - 2|=20-4 2|2x - 2|=16 |2x - 2|=8 2x-2=8 or 2x-2=-8 2x=8+2 or 2x=-8+2 2x=10 or 2x=-6 x=5 or x=-3

OpenStudy (anonymous):

ThankYou :) Your Awesomee @Nurali

OpenStudy (nurali):

My Pleasure.